OpenStudy (selena2345):

Consider The Equation below: 1/5(x+2)+2x=6x-10 Part A: Which property can be used to simplify the expression 1/5(x+2) ? Answer _______ Part B: Collect all x-terms on one side of the equation and all constants on the other side, and combine like terms. Answer _______ Part C: What is the value of x ? Answer ________

OpenStudy (anonymous):

hehe sry im not this far into math :)

OpenStudy (selena2345):

what ? @Hero

OpenStudy (selena2345):

@Hero i don't get it

OpenStudy (selena2345):

i'm confused @BraydonG

OpenStudy (selena2345):

@Hero can you help in Part B And Part C

hero (hero):

Have you finished Part A yet?

OpenStudy (selena2345):

yes

hero (hero):

What did you put for Part A?

OpenStudy (selena2345):

distributive

hero (hero):

Okay good answer.

hero (hero):

Did you attempt Part B? If so, you should post your steps.

OpenStudy (selena2345):

no

hero (hero):

In Part A, you used the distributive property, so you should have: x/5 + 2/5 + 2x = 6x - 10 For part B, you're supposed to move all the terms with x to the left side, and all the constant terms to the other side. Can you try doing that please?

OpenStudy (selena2345):

2x+6x

hero (hero):

Actually you should have x/5 + 2x - 6x = -10 - 2/5

hero (hero):

Notice all the terms with x are on the left side. All the constant terms (those without variables) on the right side.

OpenStudy (selena2345):

yes

hero (hero):

To make things easier, factor out x on the left side, and factor out the negative on the right side: (1/5 + 2 - 6)x = -(10 + 2/5)

OpenStudy (selena2345):

is this part c ? @Hero

hero (hero):

Now all you have to do is just add up the numbers for the next step

hero (hero):

Part C is to find the value of x

hero (hero):

That's what we are doing.

OpenStudy (selena2345):

ok @Hero

OpenStudy (selena2345):

@Hero

hero (hero):

Have you tried adding up the numbers yet?